# R-PINN: Recovery-type a-posteriori estimator enhanced adaptive PINN

### v1

@@ Abstract
In recent years, with the advancements in machine learning and neural networks, algorithms using physics-informed neural networks (PINNs) to solve PDEs have gained widespread applications. While these algorithms are well-suited for a wide range of equations, they often exhibit suboptimal performance when applied to equations with large local gradients, resulting in substantial localized errors. To address this issue, this paper proposes an adaptive PINN algorithm designed to improve accuracy in such cases. The core idea of the algorithm is to adaptively adjust the distribution of collocation points based on the recovery-type a-posterior error of the current numerical solution, enabling a better approximation of the true solution. This approach is inspired by the adaptive finite element method. By combining the recovery-type a-posteriori estimator, a gradient-recovery estimator commonly used in the adaptive finite element method (FEM) with PINNs, we introduce the Recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN) and compare its performance with a typical adaptive PINN algorithm, FI-PINN. Our results demonstrate that R-PINN achieves faster convergence with fewer adaptive points and significantly outperforms in the cases with multiple regions of large errors than FI-PINN. Notably, our method is a hybrid numerical approach for solving partial differential equations, integrating adaptive FEM with PINNs.
@@
Adaptive PINN , Recovery-type a-posteriori estimator , Machine learning , Adaptive FEM
@@ 1 Introduction
The Physics-informed neural networks (PINNs) algorithm is a data-driven approach for solving partial differential equations (PDEs) based on deep learning frameworks [1]. Due to its distinct methodology, which diverges significantly from traditional PDE numerical techniques, and by leveraging the superior capabilities of neural networks in handling large-scale data and optimization problems, PINNs have garnered considerable attention since their introduction and have become important tools for solving PDEs. These methods utilize automatic differentiation to compute the penalty of an equation at a set of collocation points within the domain of interest, using it as an objective loss function that is optimized through backpropagation, thereby approximating the solution of the PDE. Compared to classical PDE numerical methods such as the finite element method and finite difference method, one significant advantage of PINNs is that they are mesh-free, making them easier to implement and particularly well-suited for high-dimensional problems and irregular domains. Currently, PINNs have been applied to solve a range of PDE problems, demonstrating their strengths effectively, such as inverse problems in fluid dynamics, optics and metamaterials [2, 3], or some other forward problems [4, 5, 6].
@@
Despite the notable advantages demonstrated by PINNs in PDEs, there remains substantial room for improvement in several aspects. While PINNs have exhibited high-precision solutions across a variety of scenarios, their applicability is still limited. Notably, significant errors can arise when addressing equations such as the Allen-Cahn equation or the Burgers’ equation. Enhancements to the PINNs methodology can be pursued from multiple aspects. For instance, since the total loss function in PINNs is a weighted sum of errors from both the domain and boundary conditions, adjusting the weights of these terms may improve performance across different problems [7, 8, 9, 10]. In cases involving extensive solution domains, decomposition techniques can be employed to localize the training process, thereby enhancing optimization outcomes [11, 12, 13]. Regarding the choice of loss functions, although Mean Squared Error (MSE) is commonly used, alternative forms of loss functions may give better training results in certain circumstances [14, 10, 15]. Furthermore, strategies such as batch training and the design of specialized network architectures to deal with Dirichlet or periodic boundary conditions offer a promising direction to enhance the performance of PINNs [8, 16, 17]. In this paper, we focus on the limitations of PINNs in handling solutions with sharp local variations. This challenge, arising from singularities or high-gradient regions, is typically addressed by introducing adaptive sampling strategies for PINNs. However, existing sampling methods primarily only rely on the values of the loss function at local collocation points, without incorporating a rigorous a-posteriori error estimator. To address this issue, we propose the recovery-type a-posteriori estimator-enhanced adaptive PINN (R-PINN), which integrates the recovery-type a-posteriori estimator from classical numerical methods into PINNs. Furthermore, we introduce an adaptive sampling method, the Recovery-type estimator adaptive distribution (R-AD), to systematically implement adaptive refinement.
@@
The sampling methods for collocation points in PINNs are generally categorized into uniform and non-uniform sampling. In typical uniform sampling, points are either randomly distributed or placed at equispaced intervals, remaining fixed throughout the training process. However, this static sampling strategy has notable limitations, as it fails to capture the model’s fine local features effectively, which can lead to convergence issues in complex problems settings. To address these limitations, Lu et al. first proposed an adaptive non-uniform sampling method [18]. Since then, various adaptive sampling methods for collocation points have been developed [19, 20, 21, 22, 23, 24, 20]. Among these, the Failure-informed PINN (FI-PINN) by Gao et al. [20] is particularly notable, utilizing a truncated Gaussian model for adaptive sampling. Another significant contribution is by Wu et al. [19], who reviewed and compared a range of sampling techniques. Their work covers both uniform sampling methods, such as equispaced, random, and low-discrepancy sequences, and non-uniform methods, including residual-based adaptive refinement (RAR) and residual-based adaptive distribution (RAD), offering valuable guidance on selecting appropriate sampling strategies.
@@
While there are various adaptive PINN methods, they generally share a common approach: sampling the uniform points as background collocation points and then adding non-uniform points as the adaptive collocation points based on an a-posteriori error estimator. Most of these methods utilize the residual-type a-posteriori estimator, which calculates the error as the loss of the neural network function. For adaptive finite element methods, two primary types of estimator are widely recognized: residual-type [25, 26] and recovery-type [27, 28, 29], with the latter based on gradient recovery. Although extensive research has been conducted on adaptive PINNs, there has been limited work on gradient recovery estimators. Therefore, we introduce an adaptive PINN enhanced by the recovery-type a-posteriori estimator.
@@
The rest of this paper is organized as follows. Section 2 provides preliminary background, introducing the fundamental algorithm of Physics-informed neural networks (PINNs) and existing sampling strategies for both uniform and non-uniform purposes. In Section 3, we present the detailed algorithm of the proposed Recovery-type estimator-enhanced PINN (R-PINN). Section 4 presents our results using R-PINN on four representative singular problems, illustrating its performance and providing a comparative analysis. Finally, Section 5 summarizes our conclusions and outlines potential directions for future research.
@@ 2.1 PINNs for solving forward and inverse PDEs
Analogous to the application of basis functions for approximating true solutions in traditional numerical methods for PDEs, the Physics-informed neural networks (PINNs) algorithm utilizes deep neural networks to approximate the exact solutions of equations, primarily employing fully connected feed-forward neural networks. We denote a fully connected feed-forward neural network with $L$ layers as $\mathcal{N}^{L}:\mathbb{R}^{N_{0}}\rightarrow\mathbb{R}^{N_{k}}$, where the $k$-th layer comprises $N_{k}$ neurons. The weight matrix and bias vector of the $k$-th layer (for $1\leq k\leq L$) can be represented as $\bm{W}^{k}\in\mathbb{R}^{N_{k}\times N_{k-1}}$ and $\bm{b}^{k}\in\mathbb{R}^{N_{k}}$, respectively. The input vector is denoted as $\bm{x}\in\mathbb{R}^{N_{0}}$, and the output vector of the $k$-th layer is expressed as $\mathcal{N}^{k}(\bm{x})$, especially, $\mathcal{N}^{0}(\bm{x})=\bm{x}$. The activation function $\Phi$, applied during the feed-forward transition from the $(k-1)$-th layer to the $k$-th layer (for $1\leq k\leq L$), introduces non-linearity into the network, thereby enhancing its learning capacity. Consequently, a feed-forward neural network can be mathematically defined as
$\mathcal{N}^{k}(\bm{x})=\Phi(\bm{W}^{k}\mathcal{N}^{k-1}(\bm{x})+\bm{b}^{k}),\quad 1\leq k\leq L.$ (2.1)
Letting $\tilde{\bm{\Theta}}=\{\bm{W}^{k},\bm{b}^{k}\}\in\mathcal{V}$ be all parameters of the neural network, including weights and biases, where $\mathcal{V}$ denotes the parameter space, we can finally express the neural network as
$\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}}),$
where $\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ emphasizes the dependence of the output layer $\mathcal{N}^{L}(\bm{x})$ on $\tilde{\bm{\Theta}}$. Initial parameters are typically generated according to a specified probability distribution.
@@
$f(\bm{x};u(\bm{x}))=f\left(\bm{x};\frac{\partial u}{\partial x_{1}},\cdots,\frac{\partial u}{\partial x_{d}};\frac{\partial^{2}u}{\partial x_{1}^{2}},\cdots,\frac{\partial^{2}u}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)=0,\quad\bm{x}=(x_{1},x_{2},\cdots,x_{d})\in\Omega,$ (2.2)
@@
In solving the forward problem, where the parameter $\bm{\lambda}$ in the equation is known, PINNs approximate the true solution $u$ using a neural network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})$ [1]. During this process, the neural network is iteratively trained to update its parameters $\tilde{\bm{\Theta}}$ in order to minimize the loss function, defined as,
$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}}),$ (2.3)
@@
$\begin{split}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{f}\rvert}\sum_{\bm{x}\in\mathcal{T}_{f}}\bigg{|}&f\left(\bm{x};\frac{\partial\hat{u}}{\partial x_{1}},\cdots,\frac{\partial\hat{u}}{\partial x_{d}};\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{1}},\cdots,\frac{\partial^{2}\hat{u}}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)\bigg{|}^{2},\\ &\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}})=\frac{1}{\lvert\mathcal{T}_{b}\rvert}\sum_{\bm{x}\in\mathcal{T}_{b}}\lvert\mathcal{B}(\hat{u},\bm{x})\rvert^{2},\end{split}$ (2.4)
@@
$\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda})=\frac{1}{|\mathcal{T}_{i}|}\sum_{\bm{x}\in\mathcal{T}_{i}}|\hat{u}(\bm{x})-u(\bm{x})|^{2}.$ (2.5)
@@
$\mathcal{L}(\mathcal{T};\tilde{\bm{\Theta}},\bm{\lambda})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\tilde{\bm{\Theta}},\bm{\lambda})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\tilde{\bm{\Theta}},\bm{\lambda})+w_{i}\mathcal{L}_{i}(\mathcal{T}_{i};\tilde{\bm{\Theta}},\bm{\lambda}),$ (2.6)
with the weight of the additional term $w_{i}$. Then the unknown parameters $\bm{\lambda}$ can be trained with the neural network parameters $\tilde{\bm{\Theta}}$ by the optimization process.
@@ 2.2 Sampling methods for PINNs
Typically, the sampling strategy in PINNs involves selecting a set of collocation points by uniform sampling prior to training the neural network, with the set remaining fixed throughout the training process. Commonly employed uniform sampling methods include equispaced uniform sampling and uniformly random sampling, while other uniform sampling techniques are also available [1, 4, 30]. In Wu. et al.’s work [19], six uniform sampling methods were systematically reviewed, including equispaced uniform grid (Grid), uniformly random sampling (Random), Latin hypercube sampling (LHS) [31, 32], Halton sequence (Halton) [33], Hammersley sequence (Hammersley) [34], and Sobol sequence (Sobol) [35]. Their comparative experiments demonstrated that the Sobol method generally provides superior performance. The Sobel sequence uses a base of two to form successively finer uniform partitions of the unit interval and then reorder the coordinates in each dimension. We show the difference over three different methods in Fig. 1.
@@
Fig. 1: $900$ uniformly collocation points generated on $[0,1]^{2}$ with three different strategies.
@@
In many cases, accuracy of numerical solution can be enough through uniform sampling strategies. However, for certain numerical problems characterized by large local gradients and highly distinctive solution features, such methods may fail to adequately capture local variations. Several adaptive methods have been proposed to overcome these limitations in recent years, primarily through modifications to the sampling strategies for the collocation points in PINNs. These adaptive sampling approaches can be broadly categorized into two types [19]: the Residual-based adaptive refinement (RAR) [18] and the Residual-based adaptive distribution (RAD) [21]. RAR targets regions with large residuals, where new adaptive collocation points are introduced at regular intervals during the training process to reduce local errors (see Algorithm 1).
Algorithm 1 Residual-based adaptive refinement (RAR)
1:  Sample background collocation points $\mathcal{T}$.
2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
3:  while not reach the end do
4:     Sample a set of adaptive collocation points $\mathcal{S}_{0}$.
5:     Compute the PDE residuals of adaptive collocation points $\mathcal{S}_{0}$.
6:     $S\leftarrow m$ points with largest residuals from $\mathcal{S}_{0}$.
7:     $\mathcal{T}\leftarrow\mathcal{T}\cup\mathcal{S}$.
8:     Train the neural network for some epoches.
9:  end while
10:  Output the result.
While RAR only focuses on the area with large residuals and ignore the other place, RAD utilizes the probabilty density function $p(\bm{x})$ that has strong relationship with residual to resample all the collocation points, which can take the area both with large and small residual into care (see Algorithm 2). For example, we can take $p(\bm{x})$ proportional to the PDE residual $\epsilon(\bm{x})=|f(\bm{x};\hat{u}(\bm{x})|$,
[…]
More forms of $p(x)$ can be used here, such as it in paper[19],
$p(\bm{x})\propto\frac{\epsilon^{k}(\bm{x})}{\mathbb{E}[\epsilon^{k}(\bm{x})]}+c,$ (2.8)
where $k\geq 0,c\geq 0$ are two hyperparameters.
Algorithm 2 Residual-based adaptive distribution (RAD)
1:  Sample background collocation points $\mathcal{T}$.
2:  Train the neural network with $\mathcal{T}$ for a certain number of epoches.
3:  while not reach the end do
4:     $\mathcal{T}\leftarrow$ resampled collocation points according to the probability density function $p(\bm{x})$.
5:     Train the neural network for some epoches.
6:  end while
7:  Output the result.

### v2

@@ Abstract
In recent years, with the advancements in machine learning and neural networks, algorithms using physics-informed neural networks (PINNs) to solve PDEs have gained widespread applications. While these algorithms are well-suited for a wide range of equations, they often exhibit a suboptimal performance when applied to equations with large local gradients, resulting in substantially localized errors. To address this issue, this paper proposes an adaptive PINN algorithm designed to improve accuracy in such cases. The core idea of the algorithm is to adaptively adjust the distribution of collocation points based on the recovery-type a-posteriori error of the current numerical solution, enabling a better approximation of the true solution. This approach is inspired by the adaptive finite element method. By combining the recovery-type a-posteriori estimator, a gradient-recovery estimator commonly used in the adaptive finite element method (FEM), with PINNs, we introduce the recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN) and compare its performance with a typical adaptive sampling PINN, failure-informed PINN (FI-PINN), and a typical adaptive weighting PINN, residual-based attention in PINN (RBA-PINN) as a baseline. Our results demonstrate that R-PINN achieves faster convergence with fewer adaptively distributed points and outperforms the other two PINNs in the cases with regions of large errors.
@@
physics-informed neural networks , adaptive sampling , recovery-type estimator , gradient recovery
@@ 1 Introduction
The physics-informed neural networks (PINNs) algorithm is a data-driven approach for solving partial differential equations (PDEs) based on deep learning frameworks [26]. Due to its distinct methodology, which diverges significantly from traditional numerical techniques, and by leveraging the superior capabilities of neural networks in handling large-scale data and optimization problems, PINNs have garnered considerable attention ever since their introduction and have become important tools for solving PDEs. These methods utilize automatic differentiation to compute the residual of an equation at a set of collocation points within the domain of interest, using it as an objective loss function that is optimized through backpropagation, thereby approximating the solution of the PDE. Compared to classical PDE numerical methods such as the finite element method and finite difference method, one significant advantage of PINNs is that they are mesh-free, making them easier to implement and particularly well-suited for high-dimensional problems and irregular domains. Currently, PINNs have been applied to solve a range of PDE problems, including fluid mechanics, which has been reviewed comprehensively by Cai et al. [5], as well as fields such as optics, metamaterials and others [27, 6, 24, 38, 12].
@@
Despite the notable advantages demonstrated by PINNs in PDEs, there remains a substantial room for improvement. While PINNs have exhibited high-precision solutions across a variety of scenarios, their applicability is still limited. Notably, significant errors can arise when addressing equations such as the Allen-Cahn equation or the Burgers’ equation. Enhancements to the PINNs methodology can be pursued from multiple aspects. For instance, since the total loss function in PINNs is a weighted sum of errors from both the domain and boundary conditions, adjusting the weights of these terms may improve performance across different problems [1, 32, 19, 33, 10]. A representative example is the residual-based attention in PINN [1], which selects weighting multipliers for the loss function and has been shown to significantly enhance the performance of PINNs. In cases involving extensive solution domains, decomposition techniques can be employed to localize the training process, thereby enhancing optimization outcomes [22, 28, 16]. Regarding the choice of loss functions, although the mean squared error (MSE) is commonly used, alternative forms of loss functions may give better training results in certain circumstances [10, 25, 20]. Furthermore, strategies, such as batch training and the design of specialized network architectures to deal with Dirichlet or periodic boundary conditions, offer a promising direction to enhance the performance of PINNs [19, 17, 7].
In this paper, we focus on the limitations of PINNs in handling solutions with sharp local variations. This challenge, arising from singularities or high-gradient regions, is typically addressed by introducing adaptive sampling. However, existing sampling methods primarily rely on the PDE loss at local collocation points, which can be regarded as a residual-type estimator. Nevertheless, alternative efficient estimators remain unexplored for adaptive PINNs. Therefore, we propose the recovery-type a-posteriori estimator enhanced adaptive PINN (R-PINN), which integrates the recovery-type a-posteriori estimator from classical numerical methods into PINNs. Furthermore, we introduce an adaptive sampling method, the recovery-type estimator adaptive distribution (RecAD), which performs adaptive refinement based primarily on the recovery-type error estimator.
@@
The sampling methods for collocation points in PINNs are generally categorized into uniform and non-uniform sampling. In typical uniform sampling, points are either randomly distributed or placed at equispaced intervals, remaining fixed throughout the training process. However, this static sampling strategy has notable limitations, as it fails to capture the model’s fine local features effectively, which can lead to convergence issues in complex problems settings. To address these limitations, Lu et al. first proposed an adaptive non-uniform sampling method [18]. Since then, various adaptive sampling methods for collocation points have been developed [35, 9, 23, 8, 37, 15]. Among these, the failure-informed PINN (FI-PINN) by Gao et al. [9] is particularly notable, utilizing a truncated Gaussian model for adaptive sampling. Another significant contribution is by Wu et al. [35], who reviewed and compared a range of sampling techniques. Their work covers both uniform sampling methods, such as equispaced, random, and low-discrepancy sequences, and non-uniform methods, including residual-based adaptive refinement (RAR) and residual-based adaptive distribution (RAD), offering valuable guidance on selecting appropriate sampling strategies.
@@
While there are various adaptive PINNs methods, they generally share a common feature: sampling the uniform points as background collocation points and then adding non-uniform points as the adaptively distributed points based on an error estimator. Most of these methods utilize the residual-type a-posteriori estimator, which calculates the error as the loss of the neural network function. For adaptive finite element methods, two primary types of estimator are widely recognized: residual-type [3, 2] and recovery-type [39, 40, 41], with the latter based on gradient recovery. Although extensive research has been conducted on residual-type adaptive PINNs [35, 9], the use of gradient recovery-based estimators remains unexplored. Therefore, we introduce an adaptive PINN enhanced by the recovery-type a-posteriori estimator.
@@
The rest of this paper is organized as follows. Section 2 provides preliminary background, introducing the fundamental algorithm of physics-informed neural networks (PINNs) and existing sampling strategies for both uniform and non-uniform purposes. Section 3 presents the detailed algorithm of the proposed recovery-type estimator-enhanced PINN (R-PINN). Section 4 presents our results using R-PINN on four representative singular problems, illustrating its performance and providing a comparative analysis. Finally, Section 5 summarizes our conclusions and outlines potential directions for future research.
@@ 2.1 PINNs for solving forward and inverse PDEs
Analogous to the application of basis functions for approximating true solutions in traditional numerical methods for PDEs, the physics-informed neural networks (PINNs) algorithm utilizes deep neural networks to approximate the exact solutions of equations, primarily employing fully connected feed-forward neural networks. We denote a fully connected feed-forward neural network with $L$ layers as $\mathcal{N}^{L}:\mathbb{R}^{N_{0}}\rightarrow\mathbb{R}^{N_{k}}$, where the $k$-th layer comprises $N_{k}$ neurons. The weight matrix and bias vector of the $k$-th layer (for $1\leq k\leq L$) can be represented as $\bm{W}^{k}\in\mathbb{R}^{N_{k}\times N_{k-1}}$ and $\bm{b}^{k}\in\mathbb{R}^{N_{k}}$, respectively. The input vector is denoted as $\bm{x}\in\mathbb{R}^{N_{0}}$, and the output vector of the $k$-th layer is expressed as $\mathcal{N}^{k}(\bm{x})$, especially, $\mathcal{N}^{0}(\bm{x})=\bm{x}$. The activation function $\Phi$, applied during the feed-forward transition from the (k1)(k-1)-th layer to the $k$-th layer (for $1\leq k\leq L$), introduces non-linearity into the network, thereby enhancing its learning capacity. Consequently, a feed-forward neural network can be mathematically defined as
$\mathcal{N}^{k}(\bm{x})=\Phi(\bm{W}^{k}\mathcal{N}^{k-1}(\bm{x})+\bm{b}^{k}),\quad 1\leq k\leq L.$
Letting $\bm{\Theta}=\{\bm{W}^{k},\bm{b}^{k}\}\in\mathcal{V}$ be all parameters of the neural network, including weights and biases, where $\mathcal{V}$ denotes the parameter space, we can finally express the neural network as
$\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta}),$
where $\mathcal{N}^{L}(\bm{x};\bm{\Theta})$ emphasizes the dependence of the output layer $\mathcal{N}^{L}(\bm{x})$ on $\bm{\Theta}$. Initial parameters are typically generated according to a specified probability distribution.
@@
$f(\bm{x};u(\bm{x}))=f\left(\bm{x};\frac{\partial u}{\partial x_{1}},\cdots,\frac{\partial u}{\partial x_{d}};\frac{\partial^{2}u}{\partial x_{1}^{2}},\cdots,\frac{\partial^{2}u}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)=0,\quad\bm{x}=(x_{1},x_{2},\cdots,x_{d})\in\Omega,$
@@
In solving the forward problem, where the parameter $\bm{\lambda}$ in the equation is known, PINNs approximate the true solution $u$ using a neural network $\hat{u}_{\bm{\Theta}}(\bm{x})$ [26]. During this process, the neural network is iteratively trained to update its parameters $\bm{\Theta}$ in order to minimize the loss function, defined as,
$\mathcal{L}(\mathcal{T};\bm{\Theta})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\bm{\Theta})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\bm{\Theta}),$ (2.1)
@@
$\begin{split}\mathcal{L}_{f}(\mathcal{T}_{f};\bm{\Theta})=\frac{1}{\lvert\mathcal{T}_{f}\rvert}\sum_{\bm{x}\in\mathcal{T}_{f}}\bigg|&f\left(\bm{x};\frac{\partial\hat{u}_{\bm{\Theta}}}{\partial x_{1}},\cdots,\frac{\partial\hat{u}_{\bm{\Theta}}}{\partial x_{d}};\frac{\partial^{2}\hat{u}_{\bm{\Theta}}}{\partial x_{1}\partial x_{1}},\cdots,\frac{\partial^{2}\hat{u}_{\bm{\Theta}}}{\partial x_{1}\partial x_{d}};\cdots;\bm{\lambda}\right)\bigg|^{2},\\ &\mathcal{L}_{b}(\mathcal{T}_{b};\bm{\Theta})=\frac{1}{\lvert\mathcal{T}_{b}\rvert}\sum_{\bm{x}\in\mathcal{T}_{b}}\lvert\mathcal{B}(\hat{u}_{\bm{\Theta}},\bm{x})\rvert^{2},\end{split}$
@@
$\mathcal{L}_{i}(\mathcal{T}_{i};\bm{\Theta},\bm{\lambda})=\frac{1}{|\mathcal{T}_{i}|}\sum_{\bm{x}\in\mathcal{T}_{i}}|\hat{u}_{\bm{\Theta}}(\bm{x})-u(\bm{x})|^{2}.$
@@
$\mathcal{L}(\mathcal{T};\bm{\Theta},\bm{\lambda})=w_{f}\mathcal{L}_{f}(\mathcal{T}_{f};\bm{\Theta},\bm{\lambda})+w_{b}\mathcal{L}_{b}(\mathcal{T}_{b};\bm{\Theta},\bm{\lambda})+w_{i}\mathcal{L}_{i}(\mathcal{T}_{i};\bm{\Theta},\bm{\lambda}),$ (2.2)
with the weight of the additional term $w_{i}$. Then the unknown parameters $\bm{\lambda}$ can be trained with the neural network parameters $\bm{\Theta}$ by the optimization process.
@@ 2.2 Sampling methods for PINNs
Typically, the sampling strategy in PINNs involves selecting a set of collocation points by sampling prior to training the neural network, with the set remaining fixed throughout the training process. Common strategies for uniformly sampling collocation points include equispaced grid sampling and uniformly random sampling. Several other quasi-uniform sampling techniques have also been proposed in the literature [26, 24, 11]. In Wu et al.’s work [35], six uniform sampling methods were systematically reviewed, including equispaced uniform grid (Grid), uniformly random sampling (Random), Latin hypercube sampling (LHS) [21, 30], Halton sequence (Halton) [13], Hammersley sequence (Hammersley) [14], and Sobol sequence (Sobol) [29]. Their comparative experiments demonstrated that the Sobol method generally provides superior performance. The Sobol sequence uses a base of two to form successively finer uniform partitions of the unit interval and then reorders the coordinates in each dimension. We show the difference over three different methods in Fig. 1.
@@
Fig. 1: $900$ uniform collocation points generated on $0,1]^{2}$ with three different strategies.
@@
In many cases, accuracy of numerical solution can be good enough through uniform sampling strategies. However, for certain numerical problems characterized by large local gradients and highly distinctive solution features, such methods may fail to adequately capture the local variations. Several adaptive methods have been proposed to overcome these limitations in recent years, primarily through modifications to the sampling strategies for the collocation points in PINNs. These adaptive sampling approaches can be broadly classified into two categories:
  • The first category incrementally adds newly selected points based on an adaptive criterion, meaning that previously selected points are retained and the number of sampling points increases over time. Representative methods in this category include the residual-based adaptive refinement (RAR) [18], which identifies regions with large residuals and adds new points periodically, and the self-adaptive importance sampling (SAIS) [9], which constructs a probability density function (PDF) based on the residual to sample new points and adds them to the existing dataset.
  • The second category, in contrast, maintains a fixed total number of sampling points by periodically redistributing them based on an updated sampling criterion. A typical example is the residual-based adaptive distribution (RAD) [23], which uses a residual-informed PDF $p(\bm{x})$ to resample all collocation points. A recent study [35] shows that methods of the second category often achieve better performance. Therefore, in this work, we adopt the strategy of shifting all adaptively distributed points periodically.
The adaptive sampling methods discussed above are all based on the residual-type error estimator. However, another practical alternative, the recovery-type error estimator, has not yet been explored in PINNs-based methods. By relying on reconstructed solution fields rather than direct residuals, recovery-based estimators provide a different perspective on error characterization, which can lead to more robust and informative sampling criteria. This motivates us to develop an adaptive PINN framework based on recovery-type error estimation, along with a corresponding sampling strategy.